What is the air pressure (in bar) at a height of 75 m, if the ground level is already 1987 m above sea level? (answer to three decimal places in bars)
Answer #1

here,

as we know

pressure decreases with increasing altitude

height , h = 75 m

h2 = 1987 m

density of air , p = 1 kg/m^3

the air pressure at this height , P = P0 - p * g * ( h1 + h2)

P = 101325 - 1 * ( 1987 + 75) * 9.81

P = 8.11 * 10^4 Pa = 0.81 bar

the air pressure is 0.81 bar